OpenStudy (hatahoooochee):

Help, I'll fan and medal f(x) = x2 + 3; g(x) = Square root of quantity x minus two. Find f(g(x)).

OpenStudy (hatahoooochee):

Answers: f(g(x)) = (x2 + 3)(Square root of quantity x minus two.) f(g(x)) = Square root of quantity x minus two divided by quantity x squared plus three. f(g(x)) = x + 1 f(g(x)) = Square root of quantity x squared plus three.

OpenStudy (hatahoooochee):

@Bigbosssaint21

OpenStudy (hatahoooochee):

@griffithdylan

OpenStudy (hatahoooochee):

I got B

OpenStudy (griffithdylan):

i belive that is correct

OpenStudy (hatahoooochee):

Thank you!

OpenStudy (griffithdylan):

np